Rosas employer has instituted a flexible benefits program. Rosa will use the plan to pay for her daughters dental expenses
Answer #1

Answer:

a)

Rosa's money in Flexible Benefits Accounts=$4000

Rosa's Acutal Expense= $5000

Hence, Underestimated Expenses=Actual Expenes-Amount in Flexible Benefits Account

=$5000-$4000=$1000

Hence Cost of Underestimating Expenses=Underestimated expense*Marginal tax Bracket

=$1000*24%=$240

b)

Rosa's money in Flexible Benefits Accounts=$5000

Rosa's Acutal Expense= $4000

Hence, Overestimated Expenses=Amount in Flexible Benefits Account-Actual Expenes

=$5000-$4000=$1000

Hence Cost of Overestimating Expenses= Overestimated Expenses*(1-Marginal tax Bracket)

=$1000*(1-24%)=$760

c)

Rosa's cost of underfunding as compared with the cost of overfunding the flexible benefits account=

cost of underfunding expenses=240=240/1000*100=2.4%

cost of overestimating expenses=760=760/1000*100=7.6%

Hence we can say that while comparing the above costs underfunding will be the better option.

d)

Yes, my answer in part (c) suggest Rosa to fund the account closer to low end of her estimates.